SPECTROSCOPY OF THE NO-Ar COMPLEX IN THE VICINITY OF THE 3p ← X^{2}Π TRANSITIONS IN NITRIC OXIDE

\underline{JOE P. HARRIS}, TIMOTHY G. WRIGHT, ADRIAN M. GARDNER, ANNA ANDREJEVA, NICHOLAS A. BESLEY and OLGA V. ERSHOVA, School of Chemistry, University Park, University of Nottingham, NG7 2RD, United Kingdom; JACEK KLOS, Department of Chemistry, University of Maryland, College Park, MD 20742.

Spectra of the NO-Ar complex have been recorded using REMPI spectroscopy in the region corresponding to transitions to the 3p Rydberg states in uncomplexed NO. Additionally, spectra have been simulated using ab initio techniques with the aim of better understanding this region, as previous favourable results have been obtained using a similar approach for the 3s Rydberg state^{a}. Excellent agreement with the REMPI spectra is seen for the 3pπ state simulation. Results on the 3pσ state will be discussed, along with extension to other NO-Rare Gas complexes.

aJ. Klos, M. H. Alexander, R. Hernández-Lamoneda and T. G. Wright J. Chem. Phys., 129, 244303 (2008)